OpenStudy (anonymous):

which of the following is an odd function? a. f(x)= xcosx b.f(x)= xsinx c.f(x)= e ^cosx d.f(x)=sin^2x

OpenStudy (jamesj):

Test and see. What's the definition of an odd function?

OpenStudy (anonymous):

my answer again is letter c i dont know y.

OpenStudy (anonymous):

if a or b my answer the wiki is correct

OpenStudy (jamesj):

By definition, a function f is odd if \[ f(-x) = -f(x) \] The function in c is not odd ...

OpenStudy (anonymous):

you need to check to see if \[f(-x)=f(x)\] for even or if \[f(-x)=-f(x)\] for odd and there is no shame in trying it with numbers if the x's get confusing

OpenStudy (anonymous):

it won't be a proof, but it will give you an idea of what is going on. then you can do it with a variable

OpenStudy (jamesj):

...because \[ \cos(-x) = \cos(x) \] hence \[ e^{\cos(-x) } = e^{\cos(x)} \]

OpenStudy (anonymous):

odd x odd = even even x odd= odd even x even= even

OpenStudy (anonymous):

yun!

OpenStudy (anonymous):

ic all your answer is correct but a or b are the same right? i only need to choose 1 letter and it is letter D

OpenStudy (jamesj):

No. The function is D is not odd, because \[ \sin^2(-x) = (\sin(-x))^2 = (-\sin x)^2 = \sin^2 x \]

OpenStudy (jamesj):

hence the function in D and C are both even functions: \[ f(-x) = f(x) \]

OpenStudy (anonymous):

how could i know a and b are odd? i feel sinx and cosx are the same

OpenStudy (anonymous):

OMG its LETTER B because cos x is an even function

OpenStudy (jamesj):

No. Try and work it out by explicitly evaluating expressions.

OpenStudy (anonymous):

Geometrically, the graph of an odd function has rotational symmetry with respect to the origin, meaning that its graph remains unchanged after rotation of 180 degrees about the origin. Examples of odd functions are x, x3, sin(x), sinh(x), and erf(x).

OpenStudy (jamesj):

Yes. But stop avoiding doing the calculation. Take each of the functions in turn and evaluate f(-x). See if it equal to -f(x) or not. If it is, then f(x) is odd. That is how you answer this question.

OpenStudy (anonymous):

f(x) = -sinx is odd function so f(x) = xcosx

OpenStudy (jamesj):

f(x) = x.cos(x) is an odd function yes. Why? Because f(-x) = (-x).cos(-x) = -x.cos(x) , because cos(-x) = cos(x) = -f(x)

OpenStudy (anonymous):

its like -f(x) = f(-x) thank your mister james

OpenStudy (jamesj):

By definition, a function f(x) is odd if f(-x) = -f(x)